A bicycle and rider have a combined potential energy of 1400 J. What is the potential energy if the bicycle and rider have one fourth their mass at the same height?



1
Expert's answer
2021-10-05T10:06:32-0400

The potential energy of 1400 J can be expressed as


"E=mgh,"

the energy of reduced mass is


"E_r=\frac m4gh.\\\space\\
\frac E{E_r}=4,\\\space\\
E_r=\frac E4=350\text{ J}."